When 0.25 L of liquid nitrogen (d = 0.807 g/mL) is vaporized, what volume does the resulting gas occupy at 25 °C and 5.00 atm
1
Expert's answer
2020-05-01T14:23:51-0400

Solution.

"p \times V= \frac{m}{M}\times R\times T"

"V = \frac{m \times R \times T}{p \times M}"

"V = \frac{250 \times 0.807 \times 8.31 \times 298}{28 \times 506625} = 0.03 \ ml"

Answer:

V = 0.03 ml